# Dance-to-Music Generation with Encoder-based Textual Inversion of Diffusion Models

### v1

@@ Abstract
The harmonious integration of music with dance movements is pivotal in vividly conveying the artistic essence of dance. This alignment also significantly elevates the immersive quality of gaming experiences and animation productions. While there has been remarkable advancement in creating high-fidelity music from textual descriptions, current methodologies mainly concentrate on modulating overarching characteristics such as genre and emotional tone. They often overlook the nuanced management of temporal rhythm, which is indispensable in crafting music for dance, since it intricately aligns the musical beats with the dancers’ movements. Recognizing this gap, we propose an encoder-based textual inversion technique for augmenting text-to-music models with visual control, facilitating personalized music generation. Specifically, we develop dual-path rhythm-genre inversion to effectively integrate the rhythm and genre of a dance motion sequence into the textual space of a text-to-music model. Contrary to the classical textual inversion method, which directly updates text embeddings to reconstruct a single target object, our approach utilizes separate rhythm and genre encoders to obtain text embeddings for two pseudo-words, adapting to the varying rhythms and genres. To achieve a more accurate evaluation, we propose improved evaluation metrics for rhythm alignment. We demonstrate that our approach outperforms state-of-the-art methods across multiple evaluation metrics. Furthermore, our method seamlessly adapts to in-the-wild data and effectively integrates with the inherent text-guided generation capability of the pre-trained model. Samples are available at https://youtu.be/D7XDwtH1YwE.
Figure 1: We propose dual-path rhythm-genre inversion to incorporate rhythm and genre information from a dance motion sequence into two learnable tokens, which are then used to enhance the pre-trained text-to-music models with visual control. Through encoder-based textual inversion, our method offers a plug-and-play solution for text-to-music generation models, enabling seamless integration of visual cues. The word bank comprises learnable pseudo-words (represented as “@” for genre and “*” for rhythm) and descriptive music terms such as instruments, genres, and emotions. By combining these pseudo-words and descriptive music terms, our method allows for the generation of a diverse range of music that is synchronized with the dance.
1 Introducion
@@
If life is a sort of dance, then music is the enchanting melody that accompanies our every graceful movement. Music’s power lies in its ability to synchronize and enhance visual narratives, a trait extensively utilized in gaming and film industries. The advent of short video content has seen a surge in amateur bloggers sharing dance videos on social media, showcasing music’s integral role in this domain. Traditionally, aligning music with video has been labor-intensive, often limiting choreography to align with the music’s attributes. Our study explores using text-to-music generative models to produce music synchronized with dance videos. This approach seeks to offer an artistically and computationally efficient solution, enabling the creation of music that not only complements but also enriches the dance experience.
@@
As AI-generated content continues to advance, there has been a surge in research focusing on dance-to-music generation, aligning with contemporary trends and reflecting the broader evolution of AI applications in creative domains. Some approaches [14, 42, 9, 3, 16] employ symbolic music representation, such as MIDI and REMI, to generate music for dance videos. However, the generated music is often limited to a few classical instruments, such as piano or guitar. Furthermore, these methods often result in incoherent note sequences and simplistic melodies, deviating significantly from the characteristics of real dance music. Other approaches [54, 55, 51, 45] directly use waveforms or mel-spectrograms to generate music for dance videos, resulting in richer musical compositions. However, their audio quality is often limited, and some methods tend to overly emphasize drum beats without capturing melodic elements. Furthermore, these approaches are constrained by the genres prevalent in the dataset, making it challenging to generate diverse music suitable for dance videos. Consequently, they face difficulties when applied to real-world scenarios. Pre-trained text-to-music models have demonstrated their ability to generate high-quality and diverse music. However, most current models only provide control over global attributes of the music and lack the ability to manipulate local properties such as rhythm.
@@
To address the above-mentioned problems, we propose a novel approach for dance-to-music generation based on arbitrary pre-trained text-to-music models. Our method aims to integrate rhythm information into this model while preserving its original text-based generation capability. Given a dance video, the model generates synchronized music and allows for the manipulation of high-level features of music, ensuring a harmonious integration of music and the dance video. As illustrated in Fig. 1, our method enables dance-to-music generation in various genres (e.g., “pop”, “lock”, “house”). Additionally, it provides the ability to control high-level features of music, such as genre and emotion, using textual description. Due to our utilization of motion sequences for rhythm extraction, our method can be extended to various other physical activities (e.g., “jump rope”, “martial arts”, and “artistic gymnastics”).
@@
To achieve this goal, we seek a method to incorporate rhythm information into a pre-trained text-to-music model. This is inspired by Textual Inversion [12], which employs a pseudo-word to represent a specific concept through image reconstruction. We aim to learn a pseudo-word that represents rhythm information of arbitrary dance videos and integrate it into the vocabulary of the text encoder. However, in the text encoder, the text embeddings corresponding to an individual token are fixed, while the rhythms in different videos are variable, making it challenging to create a pseudo-word for each rhythm pattern. Therefore, we propose an encoder-based textual inversion method (inspired by [12]) that extends the text encoder with an encoder branch capable of mapping variable rhythm information to a text embedding. In contrast to prior methods [12, 53, 21, 31] using a fixed text embedding for each pseudo-word, our approach incorporates an encoder branch that enables the text embedding associated with the pseudo-word to dynamically adapt to varying input conditions. As a result, our method enables the generation of music with different rhythms based on different motion sequences in the video.
@@
  • We improve rhythm alignment evaluation metrics to align better with human perception. Experimental results demonstrate that our method achieves optimal performance on multiple evaluation metrics. Furthermore, our method seems to offer seamless adaptation to in-the-wild data.
2 Related Work
Video-aligned audio generation
@@ 2 Related Work
Video-aligned audio generation has become increasingly popular in recent years. Several approaches have explored the synthesis of sound aligned with videos [22, 33, 10, 43]. Some works utilize symbolic representation for video music generation. Audeo [41] employs a multi-stage approach, while Foley Music [14] combines graph neural networks and Transformers to generate MIDI event sequences for performance videos. RhythmicNet [42] implements a multi-stage approach for dance-to-music generation. Controllable Music Transformer (CMT) [9] designs rule-based rhythmic relationships, allowing control over timing, motion, beat, and music genre for video background music generation. DANCE2MIDI [16] constructs a dataset D2MIDI for multi-instrument MIDI and dance pairing, enabling the generation of coherent music sequences from dance videos. Video2Music [23] and V-MusProd [56] leverage multiple video features and transformer models to generate music that is synchronized with videos. However, symbolic representation-based methods overlook timbre and dynamics, limiting musical expressiveness and variation in the generated music.
@@
Recent studies have investigated the utilization of spectrograms or audio waveforms for video-to-music generation. Dance2Music-GAN (D2M-GAN) [54] is an adversarial multi-modal framework that generates complex music samples conditioned on dance videos. Zhu et al. [55] introduce a novel approach that maximizes mutual information using a conditional discrete contrastive diffusion (CDCD) loss to generate dance music. LORIS (Long-Term Rhythmic Video Soundtracker) [51] employs a latent conditional diffusion probabilistic model and context-aware conditioning encoders for synthesizing long-term conditional waveforms.  Tan et al. [45] combines a UNET-based latent diffusion model and a pre-trained VAE model to generate plausible dance music from 3D motion data and genre labels. However, most of them generate low-quality audio with significant noise, and the generated music tends to be similar, making it challenging to adapt to diverse video scenarios in the wild. In contrast, our approach, based on a pre-trained text-to-music model, is capable of generating diverse and high-quality beat-aligned dance music. V2Meow [40] utilizes a multi-stage autoregressive model to generate visually-aligned music, allowing for high-level feature control with text. However, it requires training from scratch on a large-scale dataset of O(100k), in contrast to our approach, which employs a small-scale dataset of O(1k). By leveraging encoder-based textual inversion, we drive a pre-trained text-to-music model to achieve dance-to-music generation.
@@
Text-to-music generation has made remarkable advances in recent years. Any-to-Any generation such as CoDi [46] and NExT-GPT [48] have the capability to generate across different modalities, including text-to-music generation. Several approaches like Make-an-Audio [20], AudioLDM [28], AudioLDM2 [29], and Tango [15] utilize diffusion models to generate text-guided audio encompassing speech, sounds, and music. However, these methods face limitations in terms of data quality and model scalability, resulting in lower-quality music generation. Some works such as Archisound [38], Riffusion [11], Moûsai [39], and Noise2Music [19] leverage diffusion models to achieve high-quality music generation from textual input. Furthermore, transformer-based frameworks like MusicLM [4], MUSICGEN [8], and MuseCoco [30] encode music as discrete tokens, enabling the generation of high-quality music with rich textual description. However, the above methods mainly control high-level features like genre and emotion in music generation, leaving out control over the rhythm aligned with videos. In contrast, our method utilizes learnable pseudo-words to generate video-aligned music, distinguishing it from previous approaches.
@@
While text-guided content generation has achieved impressive results, relying solely on text is insufficient for precise control over the generated content, especially when targeting specific objectives. Therefore, personalization of generative models has become a recent research focus. Personalization of generative models refers to the generation of content personalized to specific objects based on text-to-content models, leveraging prompt learning or fine-tuning techniques. DreamBooth [37] introduces a class-specific prior preservation loss for personalized image generation. LORA [18] is an efficient fine-tuning method known for achieving remarkable performance in personalized generation. Music ControlNet [49] enhances text-to-music models with composable, precise, and time-varying musical controls, without incorporating visual controls. Gal et al. [12] introduce a textual inversion (TI) method that progressively updates the embeddings of text placeholders corresponding to specific object’s visual features within a pre-trained text encoder. Inspried by TI, many variants [13, 21, 26, 47, 31, 53, 52, 5] achieve high-quality and more controllable personalized image generation. Plitsis et al. [32] directly investigate DreamBooth and TI in text-to-music generation, but this work is limited to editing audio with a reference audio. In dance-to-music generation, the incorporation of rhythm information from dance videos is necessary. Thus, as a new task of TI, we propose an encoder-based textual inversion approach that facilitates personalized music generation, allowing for the generation of music that is aligned with specific rhythmic video.
3 Method
@@
Figure 2: We employ a latent diffusion model as the generative backbone and propose an encoder-based textual inversion method. During training, we fix the prompt as “a @ music with * as the rhythm”, where “@” and “*” respectively represent the placeholders for the genre and rhythm of the input dance. Our dual-path rhythm-genre inversion optimizes the rhythm encoder and genre encoder together to generate two text embeddings that correspond to placeholders, enabling control over the genre and rhythm of the generated music in a text-to-music generation pattern. The rhythm encoder integrates traditional feature processing techniques and projector networks. Parameters $v_{i}$, $v_{i@}$, and $v_{i}$ correspond to the text embeddings of the prompt, “@”, and “*”, respectively.
@@
In this paper, we propose an encoder-based textual inversion technique, implemented as a separate network structure, that seamlessly integrates with text-to-music generation models. As shown in Fig. 2, we employ two popular text-to-music generation frameworks, namely the diffusion-based Riffusion [11] and the transformer-based MUSICGEN [8]. They are augmented with our dual-path rhythm-genre inversion, utilizing learnable pseudo-words that integrate rhythm and genre information to guide the dance-to-music generation. During training, the prompt remains fixed as “a @ music with * as the rhythm”, where “@” represents the genre description and “*” indicates the rhythm description. We obtain text embeddings for the two placeholder words using separate rhythm and genre encoders. By reconstructing the target audio, we simultaneously optimize the parameters of both encoders, facilitating their mutual enhancement. In the inference phase, flexible text descriptions can be used in addition to the fixed prompt to exert control over the high-level feature of music. The inclusion of the pseudo-word “*” representing rhythm is vital to ensure the generation of music that aligns with the dance.
3.1 Encoder-based Textual Inversion
@@
Our objective is to leverage the generative capabilities of a pre-trained text-to-music model to generate music for dance videos. Typically, a text encoder firstly tokenizes a prompt into multiple indices, each corresponding to an embedding in the corresponding embedding lookup. These text embeddings are then encoded by a text transformer to serve as conditional guidance for the generative model. For Riffusion [11], the text encoder used is the text encoder of CLIP [34], while the generative model employed is the Latent Diffusion Models (LDMs). For MUSICGEN [8], the text encoder utilized is the T5 [35], while the generative model is the Transformer decoder.
@@
One prevalent issue in most text-to-music models is their limited ability to control local attributes of music (e.g., rhythm), as these elements are challenging to describe accurately in natural language. This presents a significant challenge when generating music that aligns with dance videos, as precise control over the rhythm becomes crucial. Textual inversion is one of the personalization methods for image generation. It uses a pseudo-word (e.g., “*”), as a placeholder for a specific object and iteratively optimizes the text embedding associated with the pseudo-word. This process aims to integrate the textual description of the specific object into the vocabulary of the pre-trained text encoder, enabling targeted editing of the specific object. The optimization goal can be defined as follows:
@@
where $z\sim E(x),\epsilon\sim\mathcal{N}(0,1)$, $\epsilon_{\theta}$ denotes LDMs, $c_{\theta}$ denotes the text encoder of CLIP and $y$ denotes the prompt.
@@
To enhance the expressiveness of individual pseudo-words, we introduce an encoder branch into the text encoder. This encoder extracts features from a specific input type (e.g., human poses), and maps them to the textual space. This enables a single pseudo-word with the capacity to exhibit distinct text embeddings contingent upon the input, allowing for control over the generation of diverse content. The encoder is iteratively optimized by reconstructing the target objects. For Riffusion, the loss function of the encoder is defined as follows:
$L_{E}=\mathbb{E}_{z,x,y,\epsilon,t}[\|\epsilon-\epsilon_{\theta}(z_{t},t,t_{\theta}(x,y))\|_{2}^{2}],$ (2)
@@
$L_{\text{ce}}=\frac{1}{K}\sum_{k=1}^{K}\text{CE}(G_{\theta}(t_{\theta}(x,y))_{k},{t}_{k}),$ (3)
@@
where $t$ denotes target tokes, $x$ denotes a specific type of input, $t_{\theta}$ denotes the expanded text encoder, $G_{\theta}$ denotes the generative model and $K$ denotes the index of codebook.
@@
With a single pseudo-word, it becomes feasible to exert precise control over specific attributes of the generated content (e.g., rhythm). This control can seamlessly coexist with human text editing, resulting in enhanced flexibility. The encoder-based textual inversion is not limited to a specific task but can be applied to different areas, where the attributes are variable and challenging to describe using natural language. In our present work, we demonstrate our approach by performing dual-path rhythm-genre inversion with rhythm and genre encoders to facilitate dance-to-music generation.
3.2 Rhythm Encoder

### v2

@@ Abstract.
The seamless integration of music with dance movements is essential for communicating the artistic intent of a dance piece. This alignment also significantly improves the immersive quality of gaming experiences and animation productions. Although there has been remarkable advancement in creating high-fidelity music from textual descriptions, current methodologies mainly focus on modulating overall characteristics such as genre and emotional tone. They often overlook the nuanced management of temporal rhythm, which is indispensable in crafting music for dance, since it intricately aligns the musical beats with the dancers’ movements. Recognizing this gap, we propose an encoder-based textual inversion technique to augment text-to-music models with visual control, facilitating personalized music generation. Specifically, we develop dual-path rhythm-genre inversion to effectively integrate the rhythm and genre of a dance motion sequence into the textual space of a text-to-music model. Contrary to traditional textual inversion methods, which directly update text embeddings to reconstruct a single target object, our approach utilizes separate rhythm and genre encoders to obtain text embeddings for two pseudo-words, adapting to the varying rhythms and genres. We collect a new dataset called In-the-wild Dance Videos (InDV) and demonstrate that our approach outperforms state-of-the-art methods across multiple evaluation metrics. Furthermore, our method is able to adapt to changes in tempo and effectively integrates with the inherent text-guided generation capability of the pre-trained model. Our source code and demo videos are available at https://github.com/lsfhuihuiff/Dance-to-music_Siggraph_Asia_2024.
Dance-to-music generation; Textual inversion; Diffusion models; Pre-trained music generative models.
journalyear: 2024copyright: rightsretainedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687562isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Learning latent representationsccs: Applied computing Sound and music computing
1. Introducion
@@
If life is a sort of dance, then music is the enchanting melody that accompanies our every graceful movement. Music’s power lies in its ability to synchronize and enhance visual narratives, a trait extensively utilized in the gaming and film industries. The advent of short video content has seen a surge in amateur bloggers sharing dance videos on social media, showcasing music’s integral role in this domain. Classical approaches (Lee et al., 2005; Davis and Agrawala, 2018) align audio and video by time-warping the video. However, this process often necessitates the manual selection of suitable music and introduces variations to the original video. Additionally, copyright issues related to music can impede the widespread distribution of videos. Consequently, creating an original soundtrack for a video remains a challenging task.
@@
As AI-generated content continues to advance, there has been a surge in research focusing on dance-to-music generation, aligning with contemporary trends and reflecting the broader evolution of AI applications in creative domains. Some approaches (Gan et al., 2020; Su et al., 2021; Di et al., 2021; Aggarwal and Parikh, 2021; Han et al., 2024) employ symbolic music representation, such as MIDI and REMI, to generate music for dance videos. However, the generated music is often limited to a few classical instruments (e.g., piano, guitar). Furthermore, these methods often result in incoherent note sequences and simplistic melodies, deviating significantly from the characteristics of real dance music. Other approaches (Zhu et al., 2022a, b; Yu et al., 2023; Tan et al., 2023) directly use waveforms or mel-spectrograms to generate music for dance videos, resulting in more intricate and nuanced musical compositions. However, the generated audio quality is often suboptimal, and some tend to overemphasize percussive elements such as drum beats, failing to capture melodic nuances adequately. Moreover, these approaches are inherently limited by the musical genres prevalent in the training dataset, making it challenging to generate diverse and suitable music for dance videos across various styles. Consequently, their real-world applicability remains constrained. Pre-trained text-to-music models (Forsgren and Martiros, 2022; Copet et al., 2023) demonstrate their ability to generate high-quality and diverse music. However, most current models only provide control over global attributes of the music and cannot manipulate local properties such as rhythm.
@@
To address the above-mentioned problems, we propose a novel approach for dance-to-music generation that leverages arbitrary pre-trained text-to-music models. Our method seamlessly integrates rhythm information extracted from dance videos into these models, while preserving their original text-based generation capabilities. Given an input dance video, the model generates synchronized music that harmoniously complements the dance movements. As illustrated in Fig. 1, our method facilitates dance-to-music generation across various genres (e.g., “pop”, “lock”, “house”). Additionally, by leveraging textual descriptions, users can exert control over high-level musical attributes like genre and emotion. Notably, our approach’s reliance on motion sequences for rhythm extraction allows for its extension to diverse physical activities (e.g., “jump rope”, “artistic gymnastics”).
@@
To achieve this goal, we develop a method to incorporate rhythm information into a pre-trained text-to-music model. Inspired by Textual Inversion (Gal et al., 2023a) which employs a pseudo-word to represent a specific concept through image reconstruction, we aim to learn a pseudo-word that represents rhythm information inherent in arbitrary dance videos and integrate it into the vocabulary of the text encoder during training. However, in the text encoder, the text embeddings corresponding to an individual token are fixed, while the rhythms in different videos are variable, making it challenging to create a pseudo-word for each rhythm pattern. Therefore, in contrast to prior methods (Gal et al., 2023a; Zhang et al., 2023b; Huang et al., 2023c; Mokady et al., 2023) using a fixed text embedding for each pseudo-word, we propose an encoder-based textual inversion method that incorporates an encoder branch that enables the text embedding associated with the pseudo-word to dynamically adapt to varying input conditions. As a result, our method enables the generation of music with different rhythms based on different motion sequences in the video.
@@
  • We collect a challenging dance-music dataset entitled “In-the-wild Dance Videos” (InDV). Unlike AIST++, this dataset encompasses a wide array of dance movements within individual videos and incorporates dance genres characterized by intricate rhythmic structures, such as Chinese traditional dance.
  • Experimental results indicate that our method attains superior performance across various evaluation metrics. Additionally, our approach demonstrates robust adaptability to tempo variations and in-the-wild data.
2. Related Work
Audio-video synchronization
Some works (Lee et al., 2005; Davis and Agrawala, 2018; Sun et al., 2023) achieve audio-video synchronization through editing techniques. Among them, (Davis and Agrawala, 2018) and (Sun et al., 2023) enable alignment between audio and video and propose methods for quantifying video rhythm. Specifically, Davis and Agrawala (2018) calculate a directogram using optical flow to detect visible impacts, while Sun et al. (2023) learn visual eventfulness to capture video rhythm. While these works focus on editing methods, our approach extracts video rhythm from 2D keypoints and embeds it into a pre-trained model to generate new music that synchronizes with the videos.
@@ 2. Related Work
Generating audios synchronized with input videos (Jin et al., 2022; Qi et al., 2023; Du et al., 2023; Su et al., 2023) has become increasingly popular in recent years. Some works (Su et al., 2020; Gan et al., 2020; Su et al., 2021) use symbolic representation for video music generation. Controllable Music Transformer (CMT) (Di et al., 2021) designs rule-based rhythmic relationships, allowing control over timing, motion, beat, and music genre for background music generation. DANCE2MIDI (Han et al., 2024) constructs the D2MIDI dataset for multi-instrument MIDI and dance pairing, enabling the generation of coherent music sequences from dance videos. Video2Music (Kang et al., 2024) and V-MusProd (Zhuo et al., 2023) leverage multiple video features and transformer models to generate music that is synchronized with videos. However, symbolic representation-based methods overlook timbre and dynamics, limiting musical expressiveness and variation in the generated music.
@@
Recent studies have investigated the use of spectrograms or audio waveforms for video-to-music generation. Dance2Music-GAN (D2M-GAN) (Zhu et al., 2022a) is an adversarial multi-modal framework that generates complex music samples conditioned on dance videos. Zhu et al. (2022b) introduce a novel approach that maximizes mutual information using a conditional discrete contrastive diffusion (CDCD) loss to generate dance music. LORIS (Yu et al., 2023) employs a latent conditional diffusion probabilistic model and context-aware conditioning encoders for synthesizing long-term conditional waveforms. Tan et al. (2023) combines a UNET-based latent diffusion model and a pre-trained VAE model to generate plausible dance music from 3D motion data and genre labels. However, most of them generate low-quality audio with noticeable noise, and the generated music tends to be similar, making it challenging to adapt to diverse video scenarios in the wild. In contrast, our approach, based on a pre-trained text-to-music model, is capable of generating diverse and high-quality beat-aligned dance music. V2Meow (Su et al., 2024) utilizes a multi-stage autoregressive model to generate visually-aligned music, allowing for high-level feature control with text. However, it requires training from scratch on a large-scale O(100k) dataset, in contrast to our approach, which employs a small-scale O(1k) dataset. By leveraging encoder-based textual inversion, we drive a pre-trained text-to-music model to achieve dance-to-music generation.
@@
Text-to-music generation has made remarkable advances in recent years. Any-to-Any generation such as CoDi (Tang et al., 2023) and NExT-GPT (Wu et al., 2023c) have the capability to generate across different modalities, including text-to-music generation. LLARK (Gardner et al., 2023) combines a pre-trained music generative model with a pre-trained language model for music understanding. Several approaches like Make-an-Audio (Huang et al., 2023a), AudioLDM (Liu et al., 2023), AudioLDM2 (Liu et al., 2024), and Tango (Ghosal et al., 2023) utilize diffusion models to generate text-guided audio encompassing speech, sounds, and music. However, these methods face limitations in terms of data quality and model scalability, resulting in lower-quality music generation. Some works such as Archisound (Schneider, 2023), Riffusion (Forsgren and Martiros, 2022), Moûsai (Schneider et al., 2023), and Noise2Music (Huang et al., 2023b) leverage diffusion models to generate high quality music from textual input. Furthermore, transformer-based frameworks like MusicLM (Agostinelli et al., 2023), MUSICGEN (Copet et al., 2023), and MuseCoco (Lu et al., 2023) encode music as discrete tokens, enabling the generation of high-quality music with rich textual description. Mustango (Melechovsky et al., 2024) enhances text-to-music models with control over harmony, rhythm, and dynamics through a text format. However, these methods are limited to the single modality of music and primarily manipulate high-level features, failing to establish a connection between visual features and music. In contrast, our method employs learnable pseudo-words to achieve low-level control, enabling the generation of music that aligns with video content.
@@
While text-guided content generation has achieved impressive results, relying solely on text is insufficient for precise control over the generated content, especially when targeting specific objectives. Therefore, the personalization of generative models has become a recent research focus. Personalization of generative models refers to the generation of content personalized to specific objects based on text-to-content models, leveraging prompt learning or fine-tuning techniques. DreamBooth (Ruiz et al., 2023) introduces a class-specific prior preservation loss for personalized image generation. LORA (Hu et al., 2021) proposes an efficient fine-tuning method known for achieving remarkable performance in personalized generation. EDICT (Wallace et al., 2023) proposes an exact diffusion inversion method for image editing, which requires no model training/finetuning, prompt tuning, or extra data. Gal et al. (2023a) introduce a textual inversion (TI) method that progressively updates the embeddings of text placeholders corresponding to specific object’s visual features within a pre-trained text encoder. Inspried by TI, many variants (Gal et al., 2023b; Huang et al., 2023c; Li et al., 2023; Voynov et al., 2023; Mokady et al., 2023; Zhang et al., 2023b, a; Alaluf et al., 2023) achieve high-quality and more controllable personalized image generation. Li et al. (2024), Plitsis et al. (2024), Novack et al. (2024), Manor and Michaeli (2024) as well as Wu et al. (2023b) (MusicControlNet) explore personalized music generation using pre-trained models. However, none of these methods successfully established a connection between visual features and music. In dance-to-music generation, the incorporation of rhythm information from dance videos is necessary. Thus, as a new task of TI, we propose an encoder-based textual inversion approach that facilitates personalized music generation, allowing for the generation of music that is aligned with specific rhythmic video.
3. Method
@@
Figure 2. We employ various pre-trained music generative models as the generative backbone and propose an encoder-based textual inversion method. During training, we fix the prompt as “a @ music with * as the rhythm”, where “@” and “*” respectively represent the placeholders for the genre and rhythm of the input dance. Our dual-path rhythm-genre inversion optimizes the rhythm encoder and genre encoder together during training. Parameters $v_{i}$, $v_{i@}$, and $v_{i*}$ correspond to the text embeddings of the prompt, “@”, and “*”, respectively.
@@
In this paper, we propose an encoder-based textual inversion technique, implemented as a separate network structure, that seamlessly integrates with text-to-music generation models. They are augmented with our dual-path rhythm-genre inversion, utilizing learnable pseudo-words that integrate rhythm and genre information to guide the dance-to-music generation. During training, the prompt remains fixed as “a @ music with * as the rhythm”, where “@” represents the genre description and “*” indicates the rhythm description. We obtain text embeddings for two placeholder words using separate rhythm and genre encoders, resulting in two controllable tokens. By reconstructing the target audio, we simultaneously optimize the parameters of both encoders, facilitating their mutual enhancement. In the inference phase, flexible text descriptions can be used in addition to the fixed prompt to exert control over the high-level feature of music. The inclusion of the pseudo-word “*” representing rhythm is vital to ensure the generation of music that aligns with the dance.
3.1. Encoder-based Textual Inversion
@@
Our objective is to leverage the generative capabilities of a pre-trained text-to-music model to generate music for dance videos. Typically, a text encoder firstly tokenizes a prompt into multiple indices, each corresponding to an embedding in the corresponding embedding lookup. A text transformer then encodes these text embeddings to serve as conditional guidance for the generative model. We employ three prominent music generation frameworks as backbones: Riffusion (Forsgren and Martiros, 2022), AudioLDM (Liu et al., 2023), and MUSICGEN (Copet et al., 2023). For Riffusion and AudioLDM, the generative models employed are the Latent Diffusion Models (LDMs) (Rombach et al., 2022), with the text encoders being the text encoder of CLIP (Radford et al., 2021) and CLAP (Wu et al., 2023a), respectively. For MUSICGEN, the text encoder utilized is the T5 (Raffel et al., 2020), while the generative model is the Transformer decoder.
@@
A prevalent issue in most text-to-music models is their limited ability to control local attributes of music (e.g., rhythm), due to the difficulty of accurately describing these attributes in natural language. This limitation poses a significant challenge when generating music that aligns with dance videos, where precise control over rhythm is essential. Textual inversion (Gal et al., 2023a), a personalization method for image generation, employs a pseudo-word (e.g., “*”) as a placeholder for a specific object and iteratively optimizes the text embedding associated with the pseudo-word. This process aims to integrate the textual description of the specific object into the vocabulary of the pre-trained text encoder, thereby enabling targeted editing of the specific object. The optimization objective can be defined as follows:
@@
where $z\sim E(x),\epsilon\sim\mathcal{N}(0,1)$, $\epsilon_{\theta}$ denotes LDMs, $c_{\theta}$ denotes the text encoder of CLIP, and $y$ denotes the prompt.
@@
To enhance the expressiveness of individual pseudo-words, we introduce an encoder branch into the text encoder. This encoder extracts features from a specific input type (e.g., human poses), and maps them to the textual space. This enables a single pseudo-word with the capacity to exhibit distinct text embeddings contingent upon the input, allowing for control over the generation of diverse content. The encoder is iteratively optimized by reconstructing the target objects. For Riffusion and AudioLDM, the loss function of the encoder is defined as follows:
(2) $L_{E}=\mathbb{E}_{z,x,y,\epsilon,t}[\|\epsilon-\epsilon_{\theta}(z_{t},t,t_{\theta}(x,y))\|_{2}^{2}],$
@@
(3) $L_{\text{ce}}=\frac{1}{K}\sum_{k=1}^{K}\text{CE}(G_{\theta}(t_{\theta}(x,y))_{k},{t}_{k}),$
@@
where $t$ denotes target tokens, $x$ denotes a specific type of input, $t_{\theta}$ denotes the expanded text encoder, $G_{\theta}$ denotes the generative model and $K$ denotes the index of codebook.
@@
Utilizing a single pseudo-word enables precise control over specific attributes of the generated content (e.g., rhythm). This control can seamlessly integrate with human text editing, resulting in enhanced flexibility. The encoder-based textual inversion framework is not confined to a specific task but can be applied to different areas, where the attributes are variable and challenging to describe using natural language. In this study, we demonstrate our approach through dual-path rhythm-genre inversion, employing rhythm and genre encoders to facilitate dance-to-music generation.
3.2. Rhythm Encoder